This morning during control access, brad checked the port server connection, shielding and temperature, which are all fine.
During today's harp scan, there were continuous alarm again.
However, after the harp scan, there was no more alarm anymore. Appears
to be fixed.
A copy of this log entry has been emailed to: moffit, brads